OpenStudy (maxdrizner):

What are the intercepts of the equation -2x + 5y = 20? (-10, 0), (0, 4) (0, -10), (4, 0) (10, 0), (0, -4) (0, 10), (-4, 0)

OpenStudy (texaschic101):

To find the x intercept, sub in 0 for y -2x + 5(0) = 20 -2x = 20 x = -10....x intercept is (-10,0) To find the y intercept, sub in 0 for x -2(0) + 5y = 20 5y = 20 y = 4...y intercept is (0,4) so your intercepts are : (-10,0) and (0,4)

OpenStudy (maxdrizner):

thank you

OpenStudy (texaschic101):

no problem :)